# Symplectic Neural Flows for Modeling and Discovery

### v1

@@ Appendix D Proof of 1
where $\bar{H}$ is defined as in Equation 44, and $\widehat{H}$ as in Equation 41. This is a consequence of the fact that polynomials are continuous functions, and $\mathcal{F}$ is dense in the space of continuously differentiable scalar valued functions with respect to the $C^{1}-$topology. The proof can then be concluded by Gronwall’s inequality as in the proof of Lemma 1
We remark that in particular one can take $\varepsilon=t^{2}/N^{2}$ and get the same approximation accuracy as in Lemma 2.
@@
To conclude, we need to show that the flow map $\phi_{\bar{H},1}$ with $\bar{H}$ defined as in Equation 44 can be approximated by a suitable SympFlow. To do so, we apply the decomposition in Equation 40 to $\phi_{\bar{H},1}$. We split the time interval $[0,1]$ into $K$ substeps, leading to substeps that we denote as $\bar{\phi}_{k/K}$ associated to the integration interval $[k/K,(k+1)/K]$, with $k=0,\ldots,K-1$. We can then approximate $\bar{\phi}_{k/K}$ up to $\mathcal{O}(1/K^{2})$ with a $1-$layer SympFlow obtained by applying a Lie-Trotter splitting strategy to Equation 43. More explicitly, we can write
$\bar{\phi}_{k/K}=\phi_{\bar{H}_{1},1/K}\circ\phi_{\bar{H}_{2},1/K}+\mathcal{O}\left(\frac{1}{K^{2}}\right),$ (45)
with $\bar{H}_{1}(s,x)=\bar{V}_{\mathrm{p}}(s,p)$ and $\bar{H}_{2}(s,x)=\bar{V}_{\mathrm{q}}(s,q)$, see Corollary 1. Having $K$ of these sub-steps, we can thus conclude that there exists a SympFlow of at most $K$ layers able to approximate within $\mathcal{O}(1/K)$ the time-$1$ flow $\phi_{\bar{H},1}$. This error accumulation is due to the fact that we compose $K$ maps $\mathcal{O}(1/K^{2})$ close to the reference one. Thus, we have a global error growing like $1/K$. The $\mathcal{O}$ notation hides the multiplicative constants, which depend on the Lipschitz regularity of the flow maps.
@@
In steps 3 and 4 we showed that for each $n=0,\ldots,N-1$ there is a SympFlow of at most $K$ layers approximating within $\mathcal{O}(1/K)$ the exact flow $\phi_{\widehat{H},1}$, which is known to be $\mathcal{O}(\Delta t^{2}/N^{2})$ close to $\widetilde{\phi}_{nt/N}$, see Corollary 1. If we set $K=N^{2}$, we can thus conclude that the map $\phi_{\widetilde{H},t}$, and hence also $\phi_{H,t}$, can be approximated within $\mathcal{O}(\Delta t^{2}/N)$ by a SympFlow of at most $N^{3}$ layers. $N$ can then be selected based on the desired accuracy $\varepsilon$. This reasoning works assuming that $\Delta t$ is of moderate size, like $\Delta t=1$ we use in our experiments. Otherwise, one would have to slightly modify the reasoning.
Appendix E Description of the network architectures
In what follows, we provide details on the network architectures considered in Section 4. All the networks we compare have the same number of layers, where each layer is defined as a transformation acting on both parts, $q$ and $p$, which partition the phase-space variable $x\in\mathbb{R}^{2d}$. The number of layers chosen in the numerical experiments in Section 4 is based on SympFlow’s performance. Thus, in the unsupervised experiments, the number of layers is set to three, whereas in the supervised experiments is set to five. Additionally, in all experiments the activation function is $\sigma=\tanh$.
E.1 SympFlow
Given a SympFlow, the $i-$th layer of $\bar{\psi}_{t}$ is given by the composition map $\phi_{\mathrm{p}}^{i}\circ\phi_{\mathrm{q}}^{i}$. Each of these two composed maps depends on network weights only via a potential function, which is denoted as $V_{\mathrm{q}}^{i}$ and $V_{\mathrm{p}}^{i}$. We model both potential functions with a one-hidden layer feedforward neural network. We explicit it for $V_{\mathrm{q}}^{i}$, and the same applies to $V_{\mathrm{p}}^{i}$, possibly with different weights:
$V_{\mathrm{q}}(t,q)=\ell_{3}\circ\sigma\circ\ell_{2}\circ\sigma\circ\ell_{1}\left(\begin{bmatrix}q\\ t\end{bmatrix}\right),$
@@ E.1 SympFlow
$\ell_{1}:\mathbb{R}^{d+1}\to\mathbb{R}^{h},\,\,\ell_{2}:\mathbb{R}^{h}\to\mathbb{R}^{h},\,\,\ell_{3}:\mathbb{R}^{h}\to\mathbb{R},$
$h=10$, are three parametrized affine maps of the form $\ell_{j}(x)=A_{j}x+b_{j}$ for suitably shaped weight $A_{j}$ and bias $b_{j}$, $j=1,2,3$. For the case of the simple harmonic oscillator, where $d=1$, we thus have $30$ parameters modeling $\ell_{1}$, $110$ modeling $\ell_{2}$, and $11$ modeling $\ell_{3}$. A SympFlow of $5$ layers will hence have $5\cdot 2\cdot(30+110+11)=1510$ parameters.
E.2 MLP
for an affine map parametrized as $\ell_{i}(z)=A_{i}z+b_{i}$, $A_{i}\in\mathbb{R}^{c_{i+1}\times c_{i}}$ and $b_{i}\in\mathbb{R}^{c_{i+1}}$. For the first layer we have $c_{i}=2d+1$, since
$z=\begin{bmatrix}x\\ t\end{bmatrix},$
whereas for the last layer $c_{L+1}=2d$. All the intermediate dimensions are fixed to $c_{2}=...=c_{L}=10$. For the simple harmonic oscillator, considering an MLP of $5$ layers, we thus have that the first layer is parametrized by $40$ parameters, the intermediate three by $110$ each, and the last by $22$. This leads to a network with $40+330+22=392$ parameters.
Since composing maps as in Equation 46 does not allow to enforce the initial condition, we modify the composition and define the MLP network as
$\bar{\psi}(t,x)=x+\tanh{(t)}\cdot\varphi_{L}\circ...\circ\varphi_{1}\left(\begin{bmatrix}x\\ t\end{bmatrix}\right),$
Appendix F Additional experiments for Hénon–Heiles

### v2

@@ Appendix F Proof of Theorem 1
where $\bar{H}$ is defined as in Equation˜52. This is a consequence of the fact that polynomials are continuously differentiable functions, and $\mathcal{F}{{}_{\sigma}}$ is dense in the space of continuously differentiable scalar-valued functions with respect to the $C^{1}$-topology. Let us introduce the function $\tau:\Omega\to[0,1]$ defined as
$\tau(x)=\min\left\{1,\inf\left\{s\in[0,+\infty)\,\mid\,\phi_{\bar{H},s}(x)\in\partial\widehat{\Omega}_{\varepsilon}\right\}\right\}.$
For an $x\in\Omega$, we can then write
$\left\|\widetilde{\phi}_{t/N}(x)-\phi_{\bar{H},1}(x)\right\|_{\infty}$ $=\left\|\widetilde{\phi}_{t/N}(x)-\phi_{\widehat{H},1}(x)+\phi_{\widehat{H},1}(x)-\phi_{\bar{H},1}(x)\right\|_{\infty}$
$\leq c\frac{t^{2}}{N^{2}}+\left\|\phi_{\widehat{H},1}(x)-\phi_{\bar{H},1}(x)\right\|_{\infty}$
and hence it remains to control the last term by applying the approximation bound in (53). Let us fix $s\in[0,\tau(x)]$, a non-empty interval because of the same arguments as in the proof of Lemma 2, and consider
$\left\|\phi_{\widehat{H},s}(x)-\phi_{\bar{H},s}(x)\right\|_{\infty}$ $=\left\|\mathbf{J}\int_{0}^{s}\left(\nabla_{x}\widehat{H}(u,\phi_{\widehat{H},u}(x))-\nabla_{x}\bar{H}(u,\phi_{\bar{H},u}(x))\right)\mathrm{d}u\right\|_{\infty}$
$\leq\int_{0}^{s}\left\|\nabla_{x}\widehat{H}(u,\phi_{\widehat{H},u}(x))-\nabla_{x}\widehat{H}(u,\phi_{\bar{H},u}(x))\right\|_{\infty}\mathrm{d}u$
$+\int_{0}^{s}\left\|\nabla_{x}\widehat{H}(u,\phi_{\bar{H},u}(x))-\nabla_{x}\bar{H}(u,\phi_{\bar{H},u}(x))\right\|_{\infty}\mathrm{d}u$
$\leq\underbrace{L\int_{0}^{s}\left\|\phi_{\widehat{H},u}(x)-\phi_{\bar{H},u}(x)\right\|_{\infty}\mathrm{d}u}_{\phi_{\widehat{H},u}(x),\phi_{\bar{H},u}(x)\in\widehat{\Omega}_{\varepsilon,c}}+\underbrace{\widehat{\varepsilon}s}_{\phi_{\bar{H},u}(x)\in\widehat{\Omega}_{\varepsilon}}.$
Gronwall’s inequality then allows to obtain
$\left\|\phi_{\widehat{H},s}(x)-\phi_{\bar{H},s}(x)\right\|_{\infty}\leq\widehat{\varepsilon}s\exp(Ls)\leq\widehat{\varepsilon}\exp(L)=\varepsilon,\,\,s\in[0,\tau(x)].$
We can then show that $\tau(x)=1$ for every $x\in\Omega$. By contradiction, let us assume $0<\tau(x)<1$ for an $x\in\Omega$, i.e., $\phi_{\bar{H},\tau(x)}(x)\in\partial\widehat{\Omega}_{\varepsilon}$. By definition of $\widehat{\Omega}$, we have that $\phi_{\widehat{H},\tau(x)}(x)\in\widehat{\Omega}$, and hence
$\left\|\phi_{\widehat{H},\tau(x)}(x)-\phi_{\bar{H},\tau(x)}(x)\right\|_{\infty}\geq\min_{y\in\widehat{\Omega}}\left\|y-\phi_{\bar{H},\tau(x)}(x)\right\|_{\infty}=\varepsilon.$
This is a contradiction since we have just derived that
$\left\|\phi_{\widehat{H},\tau(x)}(x)-\phi_{\bar{H},\tau(x)}(x)\right\|_{\infty}\leq\widehat{\varepsilon}\tau(x)\exp(L\tau(x))<\widehat{\varepsilon}\exp(L)=\varepsilon.$
We conclude that $\tau(x)=1$ for every $x\in\Omega$, and hence
$\left\|\widetilde{\phi}_{t/N}(x)-\phi_{\bar{H},1}(x)\right\|_{\infty}\leq c\frac{t^{2}}{N^{2}}+\varepsilon.$
By the arbitrarity of $\varepsilon$ in the proof, we conclude that taking $\varepsilon=t^{2}/N^{2}$, it follows that there exists $\bar{H}$ realizing
$\sup_{x\in\Omega}\left\|\widetilde{\phi}_{t/N}(x)-\phi_{\bar{H},1}(x)\right\|_{\infty}=\mathcal{O}(t^{2}/N^{2}).$
@@
To conclude, we need to show that the flow map $\phi_{\bar{H},1}$ with $\bar{H}$ defined as in Equation˜52 can be approximated by a suitable SympFlow. To do so, we apply the decomposition in Equation˜47 to $\phi_{\bar{H},1}$. We split the time interval $[0,1]$ into $K$ substeps, leading to substeps that we denote as $\bar{\phi}_{k/K}$ associated to the integration interval $[k/K,(k+1)/K]$, with $k=0,\ldots,K-1$:
$\phi_{\bar{H},1}=\bar{\phi}_{(K-1)/K}\,\circ\,\cdots\circ\bar{\phi}_{2/K}\,\circ\,\bar{\phi}_{0/K},\,\,\bar{\phi}_{k/K}(x):=\phi_{\bar{H},k/K,1/K}(x).$
We can then approximate $\bar{\phi}_{k/K}$ up to $\mathcal{O}(1/K^{2})$ with a 11-layer SympFlow obtained by applying a Lie-Trotter splitting strategy to Equation˜49. More explicitly, as expanded on in Appendix˜D, we can write
$\bar{\phi}_{k/K}$ $=\bar{\phi}_{\bar{H}_{1},k/K}\circ\bar{\phi}_{\bar{H}_{2},k/K}+\mathcal{O}\left(\frac{1}{K^{2}}\right),$ (54)
$\bar{\phi}_{\bar{H}_{1},k/K}(x)$ $=\begin{bmatrix}q+\int_{k/K}^{(k+1)/K}\nabla_{p}\bar{V}_{\mathrm{p}}(s,p)\mathrm{d}s\\ p\end{bmatrix},\,\bar{\phi}_{\bar{H}_{2},k/K}(x)=\begin{bmatrix}q\\ p-\int_{k/K}^{(k+1)/K}\nabla_{q}\bar{V}_{\mathrm{q}}(s,q)\mathrm{d}s\end{bmatrix}$, (55)
with $x=(q,p)$, $\bar{H}_{1}(s,x)=\bar{V}_{\mathrm{p}}(s,p)$ and $\bar{H}_{2}(s,x)=\bar{V}_{\mathrm{q}}(s,q)$. The map $\bar{\phi}_{\bar{H}_{1},k/K}\circ\bar{\phi}_{\bar{H}_{2},k/K}$ is a SympFlow layer and, not to interrupt the flow of the proof, we show this in Lemma˜4. Having $K$ of these sub-steps, we can thus conclude that there exists a SympFlow of at most $K$ layers able to approximate within $\mathcal{O}(1/K)$ the time-$1$ flow $\phi_{\bar{H},1}$, and, on compact sets, such an approximation is uniform in the phase space variable $x$. This error accumulation is due to the fact that we compose $K$ maps $\mathcal{O}(1/K^{2})$ close to the reference one, an estimate following from the quadratic local error of Lie-Trotter, see [5]. Thus, we have a global error growing like $1/K$. The $\mathcal{O}$ notation hides the multiplicative constants, which depend on the Lipschitz regularity of the flow maps.
@@
In steps 3 and 4 we showed that for each $n=0,\ldots,N-1$ there is a SympFlow of at most $K$ layers approximating within $\mathcal{O}(1/K)$ the exact flow $\phi_{\widehat{H},1}$, which is known to be $\mathcal{O}(\Delta t^{2}/N^{2})$ close to $\widetilde{\phi}_{nt/N}$, see Corollary˜1. We recall that all these approximations can be made uniform in the phase space variable $x$ when working on compact sets. More explicitly, we have shown that
$\phi_{\widetilde{H},t}$ $=\widetilde{\phi}_{(N-1)t/N}\circ\cdots\circ\widetilde{\phi}_{0}$
$=\left(\phi_{\bar{H}_{N-1},1}+\mathcal{O}\left(\frac{\Delta t^{2}}{N^{2}}\right)\right)\circ\cdots\circ\left(\phi_{\bar{H}_{0},1}+\mathcal{O}\left(\frac{\Delta t^{2}}{N^{2}}\right)\right)$
$=\phi_{\bar{H}_{N-1},1}\circ\cdots\circ\phi_{\bar{H}_{0},1}+\mathcal{O}\left(\frac{\Delta t^{2}}{N}\right)\text{ (Step 3)},$
$\phi_{\bar{H}_{n},1}$ $=\texttt{SympFlow}_{K,n}+\mathcal{O}\left(\frac{1}{K}\right),\,n=0,...,N-1\text{ (Step 4)},$ (56)
$\phi_{\widetilde{H},t}$ $=\texttt{SympFlow}+\mathcal{O}\left(\frac{N}{K}\right)+\mathcal{O}\left(\frac{\Delta t^{2}}{N}\right),$
SympFlow $=\texttt{SympFlow}_{K,N-1}\circ\cdots\circ\texttt{SympFlow}_{K,0}.$
We remark that the notation in Equation˜56 stands for a SympFlow depending on $n$, and with at most $K$ layers. If we set $K=N^{2}$, we can thus conclude that the map $\phi_{\widetilde{H},t}$, and hence also $\phi_{H,t}$, can be approximated within $\mathcal{O}(\Delta t^{2}/N)$ by a SympFlow of at most $N^{3}$ layers. $N$ can then be selected based on the desired accuracy $\varepsilon$. This reasoning works assuming that $\Delta t$ is of moderate size, such as $\Delta t=1$, which we use in our experiments. Otherwise, one would have to slightly modify the reasoning. $\blacksquare$
Lemma 4.
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a continuously differentiable function and consider
$\bar{H}_{1}(s,x)=\bar{V}_{\mathrm{p}}(s,p)=w_{1}^{\top}\sigma(A_{1}p+b_{1}s+c_{1}),\,\bar{H}_{2}(s,x)=\bar{V}_{\mathrm{q}}(s,q)=w_{2}^{\top}\sigma(A_{2}q+b_{2}s+c_{2}),$
$\bar{H}_{1},\bar{H}_{2}\in\mathcal{F}_{\sigma}$. Assume that all the entries of $b_{1}$ and $b_{2}$ are not zero. The map $\bar{\phi}_{\bar{H}_{1},k/K}\circ\bar{\phi}_{\bar{H}_{2},k/K}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$, defined as in Equation˜55, is a layer of a SympFlow.
We note that the assumption that the entries of $b_{1}$ and $b_{2}$ never vanish is not restrictive in our case. If this were not satisfied, we could slightly perturb the entries of $b_{1}$ and $b_{2}$ and still achieve the same approximation rates. Working with zero entries would also be possible, but would result in a non-conventional SympFlow with some neurons having different activation functions from others.
Proof.
Let us focus on
$\bar{\phi}_{\bar{H}_{1},k/K}(x)=\begin{bmatrix}q+\int_{k/K}^{(k+1)/K}\nabla_{p}\bar{V}_{\mathrm{p}}(s,p)\mathrm{d}s\\ p\end{bmatrix},\,\,x=(q,p)\in\mathbb{R}^{2d}.$
We see that
$\int_{k/K}^{(k+1)/K}\nabla_{p}\bar{V}_{\mathrm{p}}(s,p)\mathrm{d}s=\int_{0}^{1/K}\nabla_{p}\bar{V}_{\mathrm{p}}\left(\frac{k}{K}+s,p\right)\mathrm{d}s,$
@@
$\bar{V}_{\mathrm{p}}\left(\frac{k}{K}+s,p\right)=w^{\top}\sigma(Ap+bs+\bar{c}),\,\,\bar{c}=c+\frac{k}{K}b,$
suppressing the weight subscripts to lighten the notation. Assume $A\in\mathbb{R}^{h\times d},\,\bar{c},b\in\mathbb{R}^{h}$. Calling $\gamma:\mathbb{R}\to\mathbb{R}$ a primitive of $\sigma$, $\gamma^{\prime}=\sigma$, we can also write
$\bar{V}_{\mathrm{p}}\left(\frac{k}{K}+s,p\right)=\sum_{i=1}^{h}\partial_{s}\left(\frac{w_{i}}{b_{i}}\gamma((Ap)_{i}+b_{i}s+\bar{c}_{i})\right)=:\partial_{s}\left(\bar{w}^{\top}\gamma(Ap+bs+\bar{c})\right),$
where $\bar{w}_{i}=w_{i}/b_{i}$, i=1,,hi=1,...,h. We can then conclude that
$\int_{k/K}^{(k+1)/K}\nabla_{p}\bar{V}_{\mathrm{p}}(s,p)\mathrm{d}s$ $=\nabla_{p}\left(\bar{w}^{\top}\gamma\left(Ap+\frac{b}{K}+\bar{c}\right)-\bar{w}^{\top}\gamma\left(Ap+\bar{c}\right)\right)$
$=\left(A^{\top}\operatorname{diag}(\bar{w})\sigma\left(Ap+\frac{b}{K}+\bar{c}\right)-A^{\top}\operatorname{diag}(\bar{w})\sigma\left(Ap+\bar{c}\right)\right).$
In summary, we see that
$\bar{\phi}_{\bar{H}_{1},k/K}(x)$ $=\begin{bmatrix}q+\int_{k/K}^{(k+1)/K}\nabla_{p}\bar{V}_{\mathrm{p}}(s,p)\mathrm{d}s\\ p\end{bmatrix}$
$=\begin{bmatrix}q+\left(A^{\top}\operatorname{diag}(\bar{w})\sigma\left(Ap+\frac{b}{K}+\bar{c}\right)-A^{\top}\operatorname{diag}(\bar{w})\sigma\left(Ap+\bar{c}\right)\right)\\ p\end{bmatrix}$
$=\begin{bmatrix}q+\nabla_{p}\tilde{V}_{\mathrm{p}}(1/K,p)-\nabla_{p}\tilde{V}_{\mathrm{p}}(0,p)\\ p\end{bmatrix}$
where
$\tilde{V}_{\mathrm{p}}(s,p)=\bar{w}^{\top}\gamma(Ap+bs+\bar{c})\in\mathcal{F}_{\gamma}.$
This allows us to conclude since $\tilde{V}_{\mathrm{p}}$ is a single-layer neural network, i.e., a particular case of an MLP. ∎
Remark 2.
Lemma˜4 also implies that SympFlows with layers based on Hamiltonians that are single hidden layer networks are universal, and one would not need to rely on more complicated MLPs from a theoretical standpoint. Practically, however, it may still be beneficial.
Appendix G Additional experiments for the Simple Harmonic Oscillator
We now include the results obtained from training SympFlow for the unsupervised experiments of the Simple Harmonic Oscillator by using the mixed training procedure. The results are in Figure˜12.
Figure 12: Mixed training procedure for the simple harmonic oscillator
Appendix H Additional experiments for Hénon–Heiles
This section collects some additional numerical experiments for the Hénon–Heiles system. Figure˜13 shows the solutions produced with the five training configurations considered in the paper. We see that the regularization is significantly improving the performance of the MLP, but not so much for the SympFlow, which already performs well just with the residual. The mixed training procedure leads to some improvements over the regularized one. Figure˜14 collects the solutions for the supervised models, whereas Figure˜15 is a reference Poincarè section for Hénon–Heiles obtained with Runge–Kutta (5,4) time integration.
@@ Appendix H Additional experiments for Hénon–Heiles
Appendix I Model comparison with varying depth
We now train and compare SympFlow and the MLP with a varying number of layers. We consider models with $4$, 88, and $16$ layers, focusing on the unsupervised experiments of the simple harmonic oscillator. For both models and all the layer numbers, we consider the training procedure with and without Hamiltonian regularization.
Table˜1 collects the average time per training epoch, and the average time per evaluation of one second of the solution for a single initial condition at inference time. These timings correspond to unsupervised experiments for the simple harmonic oscillator. We train for $10,000$ epochs, where for each we sample $700$ new collocation points to evaluate the loss function, and we also test the accuracy over $5,000$ test samples. The average training time is obtained by timing the full training loop and dividing by $10,000$. For inference, we predict the solution up to time $T=1,000$ for $N=100$ initial conditions. The total time to produce these $100$ solutions is then divided by $N\cdot T$, hence producing an average inference time per second of simulation. SympFlow is more expensive to evaluate, and further work is needed to improve its implementation efficiency. This additional cost is mostly due to the need to compute gradients for the forward pass, compared to a conventional MLP. On top of the evaluation of SympFlow, it is also evident that the Hamiltonian regularization contributes to an increased overhead cost at training time. In contrast, it has no impact on the inference time. Still, this additional cost is worth it given the considerably improved long-term reliability of the produced solution, compared to that of the MLP.
Table 1: Timings. This table collects the average training time per epoch and inference time per unit of time of four models with varying numbers of layers $L\in\{4,8,16\}$. The experiments are done on a Quadro RTX 6000 24220MiB GPU. The time is expressed in seconds ($s$), and it is measured with the Python package time.
Model Training time ($s$) Inference time ($s$)
4 8 16 4 8 16
SympFlow just residual 0.3914 0.7231 1.3931 0.2730 0.5328 1.0528
SympFlow with regularization 0.4659 0.8395 1.6222 0.2747 0.5292 1.0427
MLP just residual 0.0232 0.0326 0.05112 0.0197 0.0212 0.0226
MLP with regularization 0.0264 0.0369 0.0584 0.0199 0.0213 0.0226
In Table˜2, we instead compare the average relative $\ell^{2}$ error of the produced solution and the average relative absolute energy error for the considered models, after having applied them up to time $k\Delta t$, $\Delta t=1$, for $k=1,10,100$. These errors are defined as in the supervised experiments of the simple harmonic oscillator, see Equation˜22 for the solution error, and Equation˜23 for the error in the Hamiltonian. The models considered here are the same leading to the timings in Table˜1. We see that the SympFlow models consistently outperform the corresponding MLPs. As expected, for any number of layers, the energy behavior of SympFlow is stable, especially when compared with the one produced by MLPs. We also notice that these experiments do not display any significant improvement when the number of layers is increased. This behavior is more evident in the MLP models than for SympFlow. However, this may reflect that the deeper networks have not fully converged within $10,000$ epochs. The choice of stopping at this fixed number of epochs is made to have comparable timings of the $12$ experiments. Furthermore, focusing on the effects of regularization, we see that the performance of SympFlow tends to be slightly worse when regularized. On the other hand, the long-time energy behavior of the MLPs is consistently more stable with a regularized loss function.
Table 2: Accuracy. This table collects the relative solution and energy average errors of four models with varying numbers of layers $L\in\{4,8,16\}$. We compute the errors after applying the networks to reach the prediction time $k\Delta t$, with $\Delta t=1$ and $k\in\{1,10,100\}$. The errors correspond to an average computed over $100$ test initial conditions. The number on the side of the model name stands for the number of layers in the network. The solution errors are scaled by a factor of $10^{-3}$, whereas the energy errors by $10^{-2}$.
Model Solution error $(\cdot 10^{-3})$ Energy error $(\cdot 10^{-2})$
$\Delta t$ $10\Delta t$ $100\Delta t$ $\Delta t$ $10\Delta t$ $100\Delta t$
SympFlow with regularization, 4 $7.277$ $17.604$ $85.731$ $1.431$ $1.552$ $0.506$
SympFlow with regularization, 8 $1.983$ $17.910$ $173.982$ $0.206$ $0.580$ $0.515$
SympFlow with regularization, 16 $1.554$ $12.506$ $117.964$ $0.163$ $0.281$ $0.221$
SympFlow just residual, 4 $1.915$ $6.698$ $42.731$ $0.190$ $0.726$ $0.370$
SympFlow just residual, 8 $0.902$ $9.086$ $92.343$ $0.098$ $0.142$ $0.124$
SympFlow just residual, 16 $0.848$ $6.376$ $66.430$ $0.087$ $0.138$ $0.124$
MLP with regularization, 4 $3.292$ $14.426$ $124.304$ $0.422$ $1.786$ $12.972$
MLP with regularization, 8 $5.663$ $31.249$ $247.627$ $0.669$ $2.357$ $12.800$
MLP with regularization, 16 $8.603$ $43.347$ $276.006$ $1.102$ $6.347$ $38.693$
MLP just residual, 4 $3.986$ $20.827$ $199.441$ $0.480$ $2.158$ $16.440$
MLP just residual, 8 $5.680$ $40.154$ $325.044$ $0.759$ $4.178$ $35.408$
MLP just residual, 16 $7.398$ $52.419$ $436.470$ $0.987$ $7.287$ $49.032$
Figure˜16 provides a comparison of the loss curves while training, and over a fixed test set. We observe that SympFlow converges faster than the MLP, as it achieves better loss function values after a relatively low number of epochs. This is easier to appreciate in the models trained without regularization, where the loss functions of the two models coincide. We also observe that, for this problem, neither SympFlow nor the MLP appear to benefit from an increased number of layers, as the final training and test losses are relatively close. Still, we remark that the final loss for SympFlow is consistently smaller than that of the MLP, which supports the interest in further exploring this type of constrained architecture for these tasks. We note that the test loss only includes the residual component. This is why the regularized SympFlows have a test loss which is smaller than the training one.
(a) 4 layers
(b) 8 layers
(c) 16 layers
(d) 4 layers
(e) 8 layers
(f) 16 layers
(g) 4 layers
(h) 8 layers
(i) 16 layers
(j) 4 layers
(k) 8 layers
(l) 16 layers
Figure 16: Training and test loss: We plot the training and test loss curves for the four models trained with a varying number of layers $L\in\{4,8,16\}$.
Appendix J Comparison with Symplectic Recurrent Neural Network
We now include a comparison of the models we have considered with the Symplectic Recurrent Neural Network (SRNN) model introduced in [14]. We perform such a comparison for the supervised experiments targeting the simple harmonic oscillator.
The idea behind SRNN is to parameterize the unknown Hamiltonian energy $H:\mathbb{R}^{2d}\to\mathbb{R}$ with a scalar-valued neural network $H_{\theta}:\mathbb{R}^{2d}\to\mathbb{R}$. Then, the approximate solutions are generated with a symplectic integrator $\varphi_{\theta}^{h}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. The authors of [14] propose to minimize the loss function
$\mathcal{L}_{\text{SRNN}}(\theta)=\frac{1}{NM}\sum_{n=1}^{N}\sum_{m=1}^{M}\left\|\varphi_{\theta}^{h_{n}^{m}}\circ\cdots\circ\varphi_{\theta}^{h_{n}^{1}}(x_{n}^{0})-x_{n}^{m}\right\|_{2}^{2}.$
Since the role of this $M$ and the one used for the supervised experiments of our continuous-in-time networks is different, we focus on the case $M=1$, and $N=1500$, and compare the results corresponding to a varying noise magnitude in the training trajectories. To be able to use explicit symplectic numerical methods, the focus is on separable Hamiltonians, meaning that the neural network $H_{\theta}$ is defined based on two other functions $K_{\theta},V_{\theta}:\mathbb{R}^{d}\to\mathbb{R}$ as $H_{\theta}(q,p)=K_{\theta}(p)+V_{\theta}(q)$. We model $K_{\theta}$ and $V_{\theta}$ as MLPs with an architecture leading to a comparable number of weights to those of SympFlow. The integrator $\varphi_{\theta}^{h}$ that we use is then the Symplectic Euler method. We now compare this model with SympFlow. We remark that this separability assumption is not necessary for SympFlow.
We include the results in Table˜3. For this comparison, we reduce the $\Delta t$ from one, the usual value we consider in the paper, to $\Delta t=0.1$. This choice is because only for smaller time steps does a numerical method provide a solution which resembles the exact solution of the parametric Hamiltonian. SympFlow does not suffer from this limitation, given that $\Delta t$ is not to be interpreted as a time step, but as an upper bound on the region where the network is trained. We observe that both SRNN and SympFlow perform well overall, as a consequence of their symplectic structure. In these experiments, SympFlow achieves lower solution errors in most settings, while SRNN is mainly better at zero noise (with a few isolated wins at other noise levels).
Table 3: Comparison with SRNN. This table collects the relative solution and energy average errors of the SRNN and SympFlow models with varying numbers of noise magnitudes $\varepsilon\in\{0,0.005,0.01,0.02,0.03,0.04,0.05\}$. The training set corresponds to $N=1500$ pairs $((x_{n},h_{n}),y_{n})$, n=1,,Nn=1,...,N, of an initial condition $x_{n}$, a time step $h_{n}$ and the updated position $y_{n}\approx\phi_{H,h_{n}}(x_{n})$, where $H$ is the Hamiltonian of the simple harmonic oscillator. We sample $h_{n}\in[0,\Delta t=0.1]$ uniformly. We compute the errors after having applied the networks to reach the prediction time $k\Delta t$, with $\Delta t=0.1$ and $k\in\{1,10,100\}$. The errors correspond to an average computed over $100$ test initial conditions. The number on the side of the model name stands for noise intensity. The solution errors are scaled by a factor of $10^{-3}$, whereas the energy errors by $10^{-2}$. Lower errors for a fixed noise and prediction level are highlighted.
Model Solution error $(\cdot 10^{-3})$ Energy error $(\cdot 10^{-2})$
$\Delta t$ $10\Delta t$ $100\Delta t$ $\Delta t$ $10\Delta t$ $100\Delta t$
SympFlow, $\varepsilon=0.0$ 1.562 19.021 167.652 0.144 1.136 2.179
SRNN, $\varepsilon=0.0$ 1.347 17.166 143.446 0.119 0.892 0.786
SympFlow, $\varepsilon=0.005$ 1.708 21.037 175.532 0.178 1.632 2.401
SRNN, $\varepsilon=0.005$ 1.930 25.487 185.259 0.212 1.759 2.184
SympFlow, $\varepsilon=0.01$ 2.507 29.709 184.588 0.324 3.166 4.190
SRNN, $\varepsilon=0.01$ 2.722 33.205 184.311 0.313 3.354 5.154
SympFlow, $\varepsilon=0.02$ 2.623 28.812 184.513 0.271 2.478 3.750
SRNN, $\varepsilon=0.02$ 4.923 51.195 297.099 0.526 4.691 9.688
SympFlow, $\varepsilon=0.03$ 2.469 27.012 162.378 0.329 2.638 3.417
SRNN, $\varepsilon=0.03$ 6.155 74.184 415.789 0.737 7.302 14.453
SympFlow, $\varepsilon=0.04$ 3.493 35.479 150.782 0.456 3.882 6.734
SRNN, $\varepsilon=0.04$ 5.106 54.381 350.609 0.545 4.062 6.124
SympFlow, $\varepsilon=0.05$ 4.246 48.107 235.879 0.487 4.646 7.506
SRNN, $\varepsilon=0.05$ 13.174 132.718 499.740 2.011 18.751 31.791
Remark 3.
SympFlow and SRNN have several similarities. The main difference is that SympFlow has layers associated with the exact flows of different time-dependent Hamiltonian systems. In contrast, the SRNN has a single time-independent Hamiltonian governing the entire network, which is generated by simulating this Hamiltonian system using a symplectic integrator. To improve the robustness to noise, in [14], the authors focused on considering a larger $M$. Given the experiments in Table˜3, we see that our SympFlow has some inherent regularization in this sense. It will be worth exploring in the future different training strategies, such as those involving larger $M$ as proposed in [14].